400-N cargo rest on the inclined ladder which makes an angle of 350 with the horizontal. Find the component force on this perpendicular to the plane.


1
Expert's answer
2021-03-14T19:13:44-0400
"400\times sin 35^0=400\times 0.5736=229.44 N"